# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

### v1

@@ Appendix E Social Impact
This work introduces a practical framework that bridges realism and controllability in traffic simulation. By decoupling the pre-training and fine-tuning stages, our method enables models pre-trained on real-world datasets to be effectively adapted to physics-based simulators, preserving trajectory-level realism while improving long-horizon, closed-loop performance. This paradigm offers a viable path for data-driven approaches to transition toward physics-based simulators, facilitating more reliable closed-loop testing and training. Consequently, it contributes to the development of safer and more robust autonomous systems.

### v2

@@ Appendix E Social Impact
This work introduces a practical framework that bridges realism and controllability in traffic simulation. By decoupling the pre-training and fine-tuning stages, our method enables models pre-trained on real-world datasets to be effectively adapted to physics-based simulators, preserving trajectory-level realism and route-level controllability while improving long-horizon, closed-loop performance. This paradigm establishes a viable pathway for transitioning data-driven approaches to physics-based simulators, facilitating more reliable closed-loop testing and training. Consequently, it contributes to the development of safer and more robust autonomous systems.
@@
(a) Robustness of RIFT across diverse AV-centric traffic scenarios.
(b) Temporal stability of RIFT in closed-loop simulation.
Figure 7: Visualizations of RIFT in diverse AV-centric scenarios. (a) Robustness of RIFT across diverse AV-centric traffic scenarios. (b) Temporal stability of RIFT in closed-loop simulation. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.